OpenStudy (anonymous):

Multiply the two binomials: (2x – 2)(2x + 2) A. 2x2 – 4x + 4 B. 42x2 + 4 C. 4x - 4 D. 4x2 – 4

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] put \(a=2x,b=2\)

OpenStudy (anonymous):

4X2-4 :)